# Swimming mode determines how well mesoscale swimmers shield their odor in turbulence

### v1

@@ III Results
In summary: both swimmer types cause the odor plume to bulge vertically and shrink laterally, and also increase velocity fluctuations. As a consequence, swimmers speed up dilution of the odor downstream of their position thus shrinking the probability that a potential predator may detect them. However, pullers are more effective at olfactory shielding because they additionally dilute odor at the source, whereas pushers trap odor at the source. Note that our problem is anisotropic due to the presence of the wall and the arrangement of the swimmers within the group, as noticed by comparing odor patterns in the spanwise and wall-normal directions (Fig. 2 (a-b)). The anisotropy of the cuboid containing the swimmers is the dominant anisotropy. Indeed, when considering a cuboid with the same shape but rotated of $90$ degrees, we obtain a similar pattern as the one visualized in Fig. 2, with the $y-$ and $z-$ axis swapped (see Supplementary Fig. S2). Small differences with respect to Fig. 2 can be attributed to the anisotropy of wall turbulence (see Supplementary Fig. S4). While the shape of the group of swimmers alters the quantitative aspects of the problem, odor shielding persists for the multiple swimmer aspect ratios that we present in the supporting information (see Fig. S1 and Fig. S2).
Having examined in detail odor shielding in the base environment, we now evaluate how the shielding changes as the Reynolds numbers of the swimmers and the advecting fluid are changed. At higher swimmer Reynolds numbers, the effect of forcing by the swimmers is expected to be more significant leading to an increase in shielding efficiency with $Re_{s}$. Our numerical experiments confirm this expectation (Fig. 5 (a-d)). Interestingly, at lower swimmer $Re$ of 10 and 25, the swimmers’ dynamics have little effect on shielding the mean odor (Fig. 5 (a)). However, the fluctuations in the odor field do carry signatures of the swimming dynamics even at lower $Re$ for puller-like swimmers and consequently affect detection probabilities (Fig. 5 (b-c)). To understand how the fluid $Re$ affects shielding, we repeat our analysis for two more fluid $Re$. Our analysis indicates that the effect of odor shielding is more prominent at lower fluid Reynolds numbers (Fig. 5 (e-h)). What is critical, however, is that puller-like swimmers continue to outperform pusher-like swimmers in all three fluid $Re$ considered in our analysis.
@@
In this work, we demonstrate how the hydrodynamic fluctuations introduced by a collection of mesoscale swimmers interact with their own odor field in a turbulent flow. We focus on two prototype swimmer types, pusher-like and puller-like, and show that in both cases the hydrodynamic interactions effectively erase the group’s odor trace downstream of the swimmers. Importantly, puller-type swimmers shield odor more efficiently than pusher-like swimmers. Moreover, the effect increases at relatively low turbulence and fast swimming, providing potentially relevant mechanisms for predator avoidance. Interestingly, filter-feeding organisms like clams have been experimentally shown to frantically increase their motion upon exposure to predator cues [6]. These observations have been linked to an active response of filter feeders that increase mixing thus erasing odor more effectively [7]. Whether and how olfactory shielding may drive the evolution of different locomotion strategies or the emergence of active responses in swimming organisms remains to be tested.
@@
Supporting Information: Swimming mode determines how well mesoscale swimmers shield their odor in turbulence
@@ I Modeling Swimmer Dynamics
$F=C_{D}\rho Au_{s}^{2}/2.$ (S1)
where $C_{D}$ is the drag coefficient, $\rho$ is the density of the fluid, $A$ is the projected area of the swimmer and $u_{s}$ is the swimmer speed [28]. Following ref. [28], the values of the drag coefficient are given by
@@
Figure S1: Statistics of the mean odor field when the swimmers are arranged in a box of same width to height aspect ratio. (a) Horizontal and (b) vertical cross-sections of the mean odor field for differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers. Panels (c) and (d) show the same for puller-like swimmers. Horizontal and vertical cross-sections of the odor standard deviation differential $\Delta\sigma(\bm{x})$ for (e,f) pusher-like and (g,h) puller-like swimmers. (i) Volume $V(C)$ for both types of swimmers and for the baseline case. (j) The corresponding shielding intensities $\Delta V(C)/V_{0}(C)$. (k) Shielding intensity of the odor fluctuations $\Delta V_{\sigma}(s)/V_{\sigma 0}(s)$.
Figure S2: Statistics of the mean odor field when the swimmers are arranged in a box different aspect ratio than those presented in the main text. (a) Horizontal and (b) vertical cross-sections of the mean odor field for differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers. Panels (c) and (d) show the same for puller-like swimmers. Horizontal and vertical cross-sections of the odor standard deviation differential $\Delta\sigma(\bm{x})$ for (e,f) pusher-like and (g,h) puller-like swimmers. (i) Volume $V(C)$ for both types of swimmers and for the baseline case. (j) The corresponding shielding intensities $\Delta V(C)/V_{0}(C)$. (k) Shielding intensity of the odor fluctuations $\Delta V_{\sigma}(s)/V_{\sigma 0}(s)$.

### v2

@@ III Results
In summary: both swimmer types cause the odor plume to bulge vertically and shrink laterally, and also increase velocity fluctuations. As a consequence, swimmers speed up dilution of the odor downstream of their position thus shrinking the probability that a potential predator may detect them. However, pullers are more effective at olfactory shielding because they additionally dilute odor at the source, whereas pushers trap odor at the source. Note that our problem is anisotropic due to the presence of the wall and the arrangement of the swimmers within the group, as noticed by comparing odor patterns in the spanwise and wall-normal directions (Fig. 2 (a-b)). The anisotropy of the cuboid containing the swimmers is the dominant anisotropy. Indeed, when considering a cuboid with the same shape but rotated of $90$ degrees, we obtain a similar pattern as the one visualized in Fig. 2, with the $y$- and $z$- axis swapped (see Supplementary Fig. S3). Small differences with respect to Fig. 2 can be attributed to the anisotropy of wall turbulence (see Supplementary Fig. S6). While the shape of the group of swimmers alters the quantitative aspects of the problem, odor shielding persists for the multiple swimmer aspect ratios that we present in the supporting information [32] (see Fig. S2 and Fig. S3).
Having examined in detail the mechanism behind olfactory shielding, let us now evaluate the effect of the kinematics of the swimmers. Mesoscale swimmers such as copepods show significant positional noise within the swarm, as has been reported from field observations and experiments [6, 7]. To account for this, we model noisy kinematics of the swimmers using an Ornstein-Uhlenbeck process as explained in Section II. We consider two different values for the positional noise, a first level of noise intensity (“intermediate” $\sigma_{ms}\approx d$, where $d$ is the mean separation between the swimmers) and a vigorous noise intensity (“high” $\sigma_{ms}\approx 5d$), with the swimmers embedded in Environment 1 (Fig. 5). Comparing with Fig. 2, we see that qualitatively the mean odor differential $\Delta\bar{c}(\bm{x})$ remains similar. Puller-like swimmers persistently shield the odor, with an efficiency that decreases as the positional noise is increased. Importantly, noise does not affect the difference between swimming modes, with pusher-like swimmers robustly less efficient than puller-like swimmers, so that at high noise intensity their shielding effect vanishes (Fig. 5 (b) and S9). This suggests that the ecological advantages for puller-like swimmers are robust and preserved regardless of the intensity of positional noise.
To further corroborate the results in an experimentally observed condition, we leverage the observations reported in Refs. [6, 7, 35], where swarms of copepods were observed to maintain their position relative to the ground both in the laboratory as well as in the field. To match more closely these observations, we conduct an additional set of simulations where the swimmers’ mean position remains stationary relative to the ground, even within an incoming current of a few cm/s [3]. To this end, we change the flow Reynolds number such that the average velocity of the swimmers, relative to the surrounding flow, remains the same as in Fig. 5, but the the mean position of the swarm is fixed relative to the ground. In these simulations, the swimmers are embedded in Environment 6 and move with intermediate positional noise ($\sigma_{ms}\approx d$). Our results show that, under these conditions, both $\Delta\bar{c}(\bm{x})$ and $\Delta\sigma(\bm{x})$ are strongly shielded, with puller-like swimmers outperforming pusher-like swimmers (Fig. 6). This further illustrates that olfactory shielding by mesoscale swimmers remains significant in experimentally relevant conditions, as long as the swarm moves relative to the incoming current and whether its absolute position relative to the seafloor is fixed or not.
In this work, we demonstrate how the hydrodynamic fluctuations introduced by a collection of mesoscale swimmers interact with their own odor field in a turbulent flow. We focus on two prototype swimmer types, pusher-like and puller-like, and show that in both cases the hydrodynamic interactions effectively erase the group’s odor trace downstream of the swimmers. Importantly, puller-type swimmers shield odor more efficiently than pusher-like swimmers. Shielding is observed both in an idealized model where the relative positions of the swimmers are fixed on a grid, and in more a realistic kinematic model where the position of the swimmers is affected by noise. While noise intensity modulates the shielding efficiency, it does not affect the robust difference between swimming modalities. How the detailed statistics of the positional noise is generated by a dynamical model of swimming, and how it affects shielding efficiency is left for future work. Moreover, the effect increases at relatively low turbulence and fast swimming, providing potentially relevant mechanisms for predator avoidance and for larger and more vigorous swimmers. Further work is needed to test the ability of higher organisms to shield their own odor trace, and how it depends on the diversity of fascinating swimming modes observed in nature. Our results may be further extended to ask how swimming mode affects pheromone transport at the individual level, and in conditions that mimic pelagic lifestyle far from the seabed.
@@
Interestingly, filter-feeding organisms like clams have been experimentally shown to frantically increase their motion upon exposure to predator cues [9]. These observations have been linked to an active response of filter feeders that increase mixing thus erasing odor more effectively [2]. Whether and how olfactory shielding may drive the evolution of different locomotion strategies or the emergence of active responses in swimming organisms remains to be tested.
Data Availability
The data that support the findings of this article are openly available [15]. The direct numerical simulations are conducted using a modified, GPU-accelerated version of the code available at https://gitlab.com/vdv9265847/IBbookVdV [38].
Author Contributions
MJ and AS designed the research. FV wrote the DNS code, which MJ adapted for the present study. MJ conducted the simulations and analyzed the data. All authors contributed to the interpretation of the results. MJ and AS wrote the manuscript with input from FV.
Acknowledgements.
This research was supported by grants to AS from the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement number 101002724 RIDING) and the National Institutes of Health under award number R01DC018789. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains.
@@ Acknowledgements.
Supporting Information: Effect of swimming mode on shielding of odor traces in turbulence
@@ I Modeling Swimmer Dynamics
$F=C_{D}\rho Au_{s}^{2}/2,$ (S1)
where $C_{D}$ is the drag coefficient, $\rho$ is the density of the fluid, $A$ is the projected area of the swimmer and $u_{s}$ is the swimmer speed [13] relative to the surrounding flow. In the simulations where the swimmers are fixed to a grid, we only account for the mean relative velocity and neglect the $\sim 5$% turbulent fluctuations. For the simulations that take into account the swimmers’ kinematics, we account for the relative fluid flow, including the turbulent fluctuations. Following Ref. [13], the values of the drag coefficient are given by
@@
II Benchmarking the passive scalar simulations
To benchmark the advection-diffusion equation solver, we compare the simulation results for the odor emitted from a sphere moving through a quiescent fluid with analytical asymptotic solutions. A sphere of radius $a=1$, implemented using an immersed boundary method [39], is placed in a uniform flow. We set Reynolds number to $Re=U_{\infty}a/\nu=0.1$ and Peclet number to $Pe=U_{\infty}a/D=0.1$, where $U_{\infty}$ is the flow velocity far from the sphere. The odor field is fixed at $c(0)=c(r=a)=1$ inside the sphere. At low $Re$ and $Pe$, the leading order terms in the analytical asymptotic expansion for the concentration field near the sphere (inner solution) $c_{0}(r,\theta)$ and far from the sphere (outer solution) $C_{0}(r,\theta)$ are given by [1] {align*} c_0(r) = 1r
C_0(r,θ) = 1re^ϵr(cos(θ)-1). Here $\epsilon=Pe$ and $\theta$ is the angle from the streamwise axis. Fig. S1 shows the comparison between the numerical solution and the analytical asymptotics for $\theta=0$ and $\pi$ showing excellent agreement. Note that for $\theta=0$, both $c_{0}$ and $C_{0}$ collapses on the same curve.
Figure S1: Comparison between the numerical solution of the advection-diffusion equation and the corresponding analytical results. (a) Odor field resulting from a sphere of radius $a=1$, moving through a quiescent fluid. Here $Re=U_{\infty}a/\nu=0.1$ and $Pe=U_{\infty}a/D=0.1$. (b) Comparison of the simulation results and analytical solution along the wake centerline ($\theta=0$). (c) Comparison of the simulation results and analytical asymptotics (’inner’ and ’outer’ solutions) along the front centerline ($\theta=\pi$).
Figure S2: Statistics of the mean odor field when the swimmers are arranged in a box of same width to height aspect ratio. (a) Horizontal and (b) vertical cross-sections of the mean odor field for differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers. Panels (c) and (d) show the same for puller-like swimmers. Horizontal and vertical cross-sections of the odor standard deviation differential $\Delta\sigma(\bm{x})$ for (e,f) pusher-like and (g,h) puller-like swimmers. (i) Volume $V(C)$ for both types of swimmers and for the baseline case. (j) The corresponding shielding intensities $\Delta V(C)/V_{0}(C)$. (k) Shielding intensity of the odor fluctuations $\Delta V_{\sigma}(\theta)/V_{\sigma 0}(\theta)$.
Figure S3: Statistics of the mean odor field when the swimmers are arranged in a box with a different aspect ratio than those presented in the main text. Here, the cuboidal box containing the swimmers are vertically oriented, as opposed to the results presented in Fig. 2. Horizontal and (b) vertical cross-sections of the mean odor field for differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers. Panels (c) and (d) show the same for puller-like swimmers. Horizontal and vertical cross-sections of the odor standard deviation differential $\Delta\sigma(\bm{x})$ for (e,f) pusher-like and (g,h) puller-like swimmers. (i) Volume $V(C)$ for both types of swimmers and for the baseline case. (j) The corresponding shielding intensities $\Delta V(C)/V_{0}(C)$. (k) Shielding intensity of the odor fluctuations $\Delta V_{\sigma}(\theta)/V_{\sigma 0}(\theta)$.
@@ II Benchmarking the passive scalar simulations
Figure S5: (a) Shielding intensity of mean odor $\Delta V(C)/V_{0}(C)$ as a function of normalized threshold $C/C_{\Omega}$, (b) shielding of odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ as a function of normalized threshold $s/C_{\Omega}$, (c) normalized detection probability differential $\Delta P_{x}(x;C_{H})/P_{x}0(x;C_{H})$ as a function of the streamwise distance from the swimmers and (d) odor pdf differential $\Delta f(c;\Omega_{A}/B)$. Colors represent different swimmer Reynolds numbers (a-d) or different turbulent Reynolds numbers (e-h). Pusher-like and puller-like swimmers are represented by solid and dotted lines respectively.
@@
Figure S7: Mean odor shielding by an individual pusher-like swimmer located in the red region (Environment 1), showing that a single swimmer does not yield a significant shielding effect.
Figure S8: Effect of swimmers’ kinematics on odor shielding for an intermediate value of positional noise ($\sigma_{MS}\approx d$). A snapshot of the swimmers’ positions is shown on the top right. Vertical and horizontal cross-sections of $\Delta\bar{c}(\bm{x})$ and $\Delta\sigma(\bm{x})$ for (a) pusher-like swimmers and (b) puller-like swimmers. (c) Shielding intensities of the mean odor field $\Delta V(C)/V_{0}(C)$. (d) Shielding intensity of the odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ for pusher-like and puller-like swimmers.
Figure S9: Effect of swimmers’ kinematics on odor shielding for a high value of positional noise ($\sigma_{MS}\approx 5d$). A snapshot of the swimmers’ positions is shown on the top right. Vertical and horizontal cross-sections of $\Delta\bar{c}(\bm{x})$ and $\Delta\sigma(\bm{x})$ for (a) pusher-like swimmers and (b) puller-like swimmers. (c) Shielding intensities of the mean odor field $\Delta V(C)/V_{0}(C)$. (d) Shielding intensity of the odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ for pusher-like and puller-like swimmers.